Operation Summary for August 24, 2011 - August 26, 2011 |
Wednesday August 24 The day shift began with operators conducting Tevatron (TEV) orbit checks, with Operations monitoring the Recycler stash (327.0E10), with the Antiproton Source (Pbar) stacking (69.8E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store. During the day shift, at 8:05 AM, the Cryo System Coordinator reported that the TEV C3 sector wet engine required adjustments; this work was completed 11 minutes later. Operators reported at 8:19 AM that the Pelletron was on. At 9:26 AM, the NS1 LCW system tripped off for the first of three times in three hours; it reset. At 9:55 AM, a Booster RF station (BRF15) tripped off on a ground fault indication and wouldn’t reset; a Booster RF person investigated and reported that the ground fault was in the tunnel and bypassed the station. Operations put the TEV into a proton only store at 11:02 AM. At 11:33 AM, BRF16 tripped off and would not reset. Operations put the TEV into a 3x3 study store at 12:02 PM and terminated it at 1:28 PM. Booster RF technicians brought BRF15 back online. Operations put the TEV into a 3x3 study store at 2:47 PM. Operations established beam to the Meson FTBFT-1015experiment at 2:59 PM. Operators reported a MI-10 vacuum burst at 3:06 PM. During the evening shift, operators reported at 5:21 PM that the Pelletron tripped off; it reset. Operators terminated the 3x3 store and then established a 3x0 proton only store. At 5:57 PM, the TEV quenched at TEV sector E1. At 6:07 PM, LRF3 tripped off; Linac personnel had it back on line 30 minutes later. Operators put the TEV through a dry squeeze at 7:38 PM. Operators reported at 7:47 PM that a Linac quadrupole magnet power supply (L:Q73) was producing large current spikes; operators called in a Linac technician to investigate. Operations turned off beam to T-1015 at 8:07 PM. Operations put the TEV into shot setup at 8:41 PM. A Linac tech reported at 9:30 PM that he replaced the control chassis power supply for L:Q73. Operations established store 9028 at 9:45 PM with an average initial luminosity (AIL) of 291.2E30. At 10:50 PM, operators reported that the TEV luminosity was jumping around; they contacted a TEV person to help find the problem. The TEV luminosity stopped jumping around at 11:51 PM. Operators reported reduced intensity out of the Booster during the shift due to the loss of two RF stations, BRF6 and BRF16 Thursday August 25 The midnight shift began with Operations monitoring store 9028 (182.53E30) and the stash (201.5E10), with Pbar stacking (8.6E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operators reported at 1:34 AM that they found a TEV separator (C:B4SVP) had caused the luminosity to jump around during the end of the last shift. The Neutron Therapy Facility (NTF) began tuning up at 7:15 AM in preparation for treating two patients later in the morning. Operators reported short downtimes due to a MI power supply (I:VT703), BRF14, and a Preaccelerator I- Source vacuum valve. During the day shift, at 8:09 AM, LRF3 tripped off; it reset. Linac techs reported at 9:13 AM that they were taking down the H- Source for maintenance; the H- Source was the backup source. Operations terminated store 9028 at 10:02 AM. Operations established a 3x3 study store at 11:18 AM. A Linac Klystron RF station (KRF1) tripped off at 11:30 AM for the first of three times in seven minutes; it reset. Operators had Booster prepared for access at 12:28 PM for RF work. At 12:43 PM, the Pelletron tripped off and suffered a full discharge; Pelletron personnel put it into recovery mode, but it continued to trip off. At 1:01 PM, the TEV 3x3 store aborted when a contactor for T:SF opened. Operators put the TEV through a dry squeeze at 2:46 PM. During the evening shift, operators reported that Booster came out of access at 4:11 PM and work on the West Anode Power Supply (WAPS) and it glycol skid continued. Pelletron personnel made a one hour access to MI-31 at 5:19 PM. Booster came back online at 5:34 PM; Operations established beam to NuMI and Pbar at 5:38 PM. Operators reported at 5:44 PM that the MiniBooNE horn charging power supply kept tripping off; they contacted MiniBooNE personnel. Operations established beam to T-1015 around 6 PM. Operations put the TEV into a proton only store at 6:22 PM. A Pelletron person reported at 6:22 PM that there was a problem with the fiber optic system inside the Pelletron and they needed an access. Operations put the TEV into shot setup at 6:45 PM. Operations established beam to MiniBooNE around 7:04 PM. Operations established store 9032 at 7:58 PM with an AIL of 199.53E30. Operations turned off beam to T-1015 at 8:29 PM. At 10:52 PM, operators reported that a transfer of antiprotons to the Recycler was lost when a Front End (FE) stopped polling; operators rebooted the FE. Friday August 26 The midnight shift began with the Pelletron off and being prepared for access, with Operations monitoring store 9032 (105.98E30) and the stash (160.7E10), with Pbar stacking (70.3E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 12:24 AM, the WAPS tripped off for the first of five times throughout the shift; it reset. Operators reported at 3:45 AM that the Pelletron was ready for access. At 4:31 AM, LRF3 tripped off and it took operators 45 minutes before they could get the station to stay on. Pbar resume stacking at 5:16 AM. LRF3 tripped off again at 7:10 AM; a Linac person investigated. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow TEV personnel another 3x3 study around noon today. She will also allow Linac and Booster to go down and make accesses if necessary for repairs. MTA requested beam for this afternoon, but the Run Coordinator said they would have to wait and see how the other work went. |
|
More Information Comments and Suggestions |
| last modified 08/26/2011 email Fermilab |
| Security, Privacy, Legal | |